A massless spring with force constant k launches
a ball of mass m . In order for the ball to reach a
speed u , by what displacement s should the
spring be compressed?


please give process

Grade:9

1 Answers

Sunil Kumar FP
10 years ago
suppose the spring is ompressed such that it reaches at height h at some intermediate level.
at that point applying law of conservation of energy
ke + pe=initial energy stored in the spring
1/2mv^2+1/2mv^2=ks^2
mv^2=ks^2
s=v under root m/k

thanks and regards
sunil kr
askIItian faculty